OpenStudy (anonymous):

-8x -24 = -26 2over3 I don't know how to make the fraction sign in here so I hope this isn't too confusing.

OpenStudy (saifoo.khan):

is it -26 2/3 ?? mixed fraction?

OpenStudy (anonymous):

Yes

OpenStudy (blacksteel):

Okay, we get -8x = -2 2/3, or -8x = - 8/3, or x = 1/3.

OpenStudy (saifoo.khan):

ahh, my bad 1/3